# ContextGen: Contextual Layout Anchoring for Identity-Consistent Multi-Instance Generation

### v2

@@ 4.2 Baselines
We compare our method against a comprehensive set of state-of-the-art baselines across relevant domains. For layout-to-image generation, we include pioneering works such as LAMIC (chen2025lamic) and MS-Diffusion (wang2025msdiffusion). To evaluate spatial control, we also benchmark against CreatiLayout (zhang2024creatilayout), EliGen (zhang2025eligenentitylevelcontrolledimage), MIGC (zhou2024migc), 3DIS (zhou20243dis), InstanceDiffusion (wang2024instancediffusion), and GLIGEN (li2023gligen). In the domain of subject-driven generation, we benchmark against OmniGen2 (wu2025omnigen2), DreamO (mou2025dreamo), UNO (wu2025less), XVerse (chen2025xverse), and MIP-Adapter (huang2024resolvingmulticonditionconfusionfinetuningfree) to specifically assess identity preservation. For a cutting-edge benchmark, we highlight the latest proprietary models, including Nano Banana (formally named Gemini 2.5 Flash Image, Google’s latest multimodal model) (gemini25flash) and GPT-4o-Image (SimaAddendumTG), as well as leading open-source models like Qwen-Image-Edit (wu2025qwenimagetechnicalreport) and FLUX.1-Kontext (labs2025flux1kontextflowmatching).
@@ 4.3 Comparison
Table 2: Quantitative results on COCO-MIG and LayoutSAM-Eval. Image-guided methods () use our pre-generated images by FLUX.1-Dev (flux2024).
@@ 4.4 Ablation Study
Prior work (zhou2025dreamrenderer) has demonstrated that MID-19 blocks have the most significant influence on instance-specific attributes. In alignment with this finding, our experiments confirm that applying the ICA mechanism selectively to the MID-19 blocks yields the highest average score of 64.66 and the best IDS score of 32.72. Furthermore, the results in Tab. 3 suggest some potential functional focus across the FLUX-DiT layers, with a detailed analysis provided in Sec. B.6.
@@
To mitigate the model’s tendency to rigidly copy layout images while neglecting instance adaptation (e.g., posture, lighting), we employ Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage), utilizing target images as preferred samples and layout images as less preferred. We first conduct an ablation study on LoRA Rank 256 to determine the optimal $\beta$ coefficient (results summarized in Tab. 4). The optimal $\beta$ (1000) was subsequently applied to our final DPO fine-tuning, which uses the higher capacity LoRA Rank 512. The qualitative validation for the final model is provided in Sec. B.3.3.
@@ A.1 Base Model Selection
In our framework, the selection of the base diffusion model is crucial for achieving high-fidelity multi-instance generation. We evaluated three variants of the FLUX family of models as potential backbones: FLUX.1-Dev (a general image generation model) (flux2024), FLUX.1-Fill (a local inpainting model) (flux1_fill_dev), and FLUX.1-Kontext (an editing model) (labs2025flux1kontextflowmatching). While existing multi-subject-driven generation methods without layout control (wu2025less; hu2025positionicunifiedpositionidentity) that do not rely on attention masks have successfully utilized FLUX.1-Dev, our experiments showed a significant limitation: without additional fine-tuning, FLUX.1-Dev failed to produce coherent images when an attention mask was applied. In contrast, both FLUX.1-Fill and FLUX.1-Kontext demonstrated the ability to generate images correctly with the attention mask. Among these two, FLUX.1-Kontext exhibited a noticeably superior capacity for identity preservation. Therefore, we chose FLUX.1-Kontext as the foundational model for our framework, leveraging its robust out-of-the-box performance with attention masking and its strong identity preservation capabilities.
@@ A.2 Details of Compositing Layout Image
Our Contextual Layout Anchoring (CLA) mechanism relies on a meticulously constructed composite layout image to achieve robust spatial control. This process involves two key steps: determining the optimal composition order for all instances and then precisely placing each instance onto the canvas.
A correct composition order is crucial for multi-instance synthesis, especially when handling occlusions and complex overlaps. We propose a dynamic sorting algorithm, Instance Layering Prioritization, which first handles explicit containment relationships by prioritizing instances whose masks are completely contained within another’s. For all other candidate instances, we use a hybrid priority scoring system to simulate the natural layering of objects. We utilize a pre-processing step to obtain each instance’s precise effective area (ravi2024sam2). The priority score $P_{i}$ is calculated as:
$P_{i}=\alpha\cdot\mathcal{A}(\text{instance}_{i})+\beta\cdot\left(1-\sum_{j\neq i}\text{IoU}(\text{instance}_{i},\text{instance}_{j})\right)+\lambda\cdot\text{RandomFactor}$
@@ B.1 Experimental Details on LAMICBench++
As shown in Tab. 1, we benchmark our method against several strong baselines, including single-image-editing models (wu2025qwenimagetechnicalreport; labs2025flux1kontextflowmatching) and closed-source commercial models (openai2024gpt4ocard; gemini25flash). Since these two categories of models require different input modalities and instruction methods, we prepared distinct inputs for a fair evaluation:
@@ B.2 Ablation on Position Indexing
As specified in Sec. 3.2, we extend the position indexing strategy from prior work (wu2025less), which uses non-overlapping indices for explicit spatial delineation of objects, crucial for multi-instance distinction. The ablation results on LAMICBench++ are presented in Tab. 5. Our findings validate the importance of the positional encoding components:
@@ B.6 Analysis on Layer-wise ICA Ablation
  • F-19: Strategy to Minimize Modality Interference. Prior work CreatiLayout (zhang2024creatilayout) extending the F-19 blocks has shown a powerful control in both spatial and attribute. And as stated this work, no further modification is made to the other blocks (Single Stream Blocks, i.e., Mid-19 and Back-19 blocks) in order to reduce the potential interference among modalities (i.e., text, layout, and image). It does make sense in considering the Front-19 blocks contains huge amount of parameters than the others. Adding ICA to these blocks is fairly enough, while applying ICA to more blocks lead to some side effect like identity distortions. That aligned with our observation that $F+M+B$ and $F+B$ strategies underperform others.
  • M-19: Critical Role in Attribute Binding. Prior work DreamRenderer (zhou2025dreamrenderer) demonstrates through ablation study that Mid-19 blocks play a relatively more critical role in attribute binding. Our ablation study validate the finding, for applying ICA to Mid-19 blocks yields the best performance in our task on benchmark score.
@@ B.7 MS-FLUX: A Controlled Case Study of Strategy Failure on Strong Backbone
Previous image generation methods based on older diffusion models (e.g., Stable Diffusion) have shown considerable improvements over their base architectures. For instance, MS-Diffusion (wang2025msdiffusion), built on SDXL, achieved image-guided layout control via its Grounding Resampler mechanism based on SD-Unet cross attention architecture. This Resampler fuses reference image content with explicit spatial and phrase information into multi-modal prompt tokens to guide generation.
@@
  • Resampler Initialization and Input Preparation. The Grounding Resampler is initialized within custom LightningModule to match the dimensions of the FLUX Transformer. Official implementation of Grounding Resampler is available at https://github.com/MS-Diffusion/MS-Diffusion/blob/main/msdiffusion/models/projection.py. During the training step, image, layout, and phrase information are prepared.
@@ C.1 Brief Survey and Discussion on Existing Multi-Instance Datasets
Various approaches have been adopted to construct multi-instance datasets. Prior works like OmniGen series (wu2025omnigen2; xiao2024omnigen) and MS-Diffusion (wang2025msdiffusion) use a cross-verification method to match and pair group images and corresponding individual images from web images or video frames. And some methods begin to leverage synthetic data for training. For example, UNO (wu2025less) employs a co-evolutionary data generation and training process using the model itself, and XVerse (chen2025xverse) add data generated by FLUX.1-Dev as supplementary to high-aesthetic-quality images.
@@ C.2 More Details of IMIG-100K Dataset
We have designed multiple prompt generation templates to include more scenes and styles. These templates are randomly combined and fed to LLMs maximize the diversity of the resulting image generation prompts. We also use different LLMs (i.e. DeepSeek v3 (deepseekai2025deepseekv3technicalreport), Gemini 2.5 Flash (gemini25flash) and GPT-4o (openai2024gpt4ocard)) for lexical and stylistic variations. Besides the realistic style (natural style), as illustrated in Fig. 12, we also include extra styles like fantasy and anime, which account for more user scenarios. The proportion of natural scenes remains high, accounting for approximately 85% to 90% of the total data.
@@
Every instance’s is meticulously annotated with detailed spatial information, including the instance phrase, its valid area’s bounding box and mask, and its corresponding position within the ground-truth image. The third subset of IMIG-100K—the flexible composite part—features more detailed annotations. Recognizing that many user scenarios only provide a facial avatar instead of a half or full-body image, we use face-related tools (facexlib; guo2021sample) to restore and standardize the faces in reference images into aligned facial images, as shown in Fig. 14. Face pairs in the ground-truth images (composited images) and corresponding reference images are strictly filtered by face quality and similarity. Crucially, the positions of valid faces are annotated (indicated by the green round bounding boxes in Fig. 14) across composited images, corresponding reference images and aligned face images. This enables a precise position and scale alignment when positioning and sizing reference instances in canvas in training process, instead of relying on a very rough bounding box of the entire instance.
@@
Our dataset has undergone strict quality assessment (kirstain2023pickapicopendatasetuser; Boutros_2023_CVPR) for both reference images and ground-truth images. Additional samples illustrating the quality and annotation detail are provided in Figs. 13 and 14.
@@ D.2 More Results on COCO-MIG
Full quantitative result on COCO-MIG benchmark is shown in Tab. 8. Our method establishes a new state-of-the-art on all key metrics, achieving the highest average Success Rate (33.12%) and mIoU (65.12%) among all compared methods. Notably, our performance advantage is most pronounced in complex, high-instance-count scenarios. For $L_{4}$, $L_{5}$, and $L_{6}$ levels, our method significantly outperforms all baselines with a Success Rate of 28.12%, 23.12%, and 24.38%, respectively. This demonstrates the robust scalability of our hierarchical architecture to maintain both layout and identity control in intricate scenes. While some competitors show slightly higher scores on individual metrics like Global Clip (CreatiLayout (wu2025less)) or Local Clip (InstanceDiffusion (wang2024instancediffusion)), our approach achieves a superior overall balance. The consistently high mIoU scores across all complexity levels and our leading Instance Success Rate on the most challenging cases further validate our model’s ability to master both precise instance placement and high-fidelity attribute preservation.
More qualitative results on COCO-MIG are shown in Fig. 17. Our method demonstrates a clear qualitative advantage over existing models on the COCO-MIG benchmark. In the first example, other methods fail to correctly generate the blue vase, with issues ranging from incorrect position to instance merging. Our model, in contrast, precisely renders the vase as intended. Similarly, for the ”green potted plant”, our approach correctly applies the “green” attribute to the pot, whereas competitors fail to do so. This highlights our superior ability to handle holistic subject identity. Furthermore, while some baselines correctly generate the requested objects in the third and fourth examples, their outputs often lack aesthetic harmony and visual coherence. Our method consistently produces images that are not only accurate but also visually pleasing and well-composed. These results underscore two key advantages of our framework: (1) Compared to other image-guided methods like MS-Diffusion (wang2025msdiffusion), our approach offers significant superiority in layout control (2) Our dedicated identity preservation mechanism provides more robust and reliable subject fidelity than the attribute-based control of text-guided methods, particularly in intricate, multi-instance scenes.
@@ Appendix E Limitations and Future Work
While our framework demonstrates state-of-the-art performance in multi-instance generation, it is not without limitations. A primary challenge stems from our model’s strong emphasis on identity preservation. When inconsistencies exist between the provided reference images or between the images and the text prompt, our model tends to prioritize maintaining the identities of the reference subjects. This can sometimes lead to a lack of flexibility in adjusting attributes such as lighting, color, or pose, which may compromise the overall visual harmony and text-image consistency of the final output. This trade-off between identity fidelity and contextual flexibility represents an important area for future research. In the future, we plan to explore more dynamic attention mechanisms that can better balance these competing demands, allowing for more flexible style and attribute transfer while preserving core subject identities.

### v3

@@ 4.2 Baselines
We compare our method against a comprehensive set of state-of-the-art baselines across relevant domains. For layout-to-image generation, we include pioneering works such as LAMIC (Chen et al., 2025b) and MS-Diffusion (Wang et al., 2025). To evaluate spatial control, we also benchmark against CreatiLayout (Zhang et al., 2024), EliGen (Zhang et al., 2025a), MIGC (Zhou et al., 2024b), 3DIS (Zhou et al., 2024c), InstanceDiffusion (Wang et al., 2024), and GLIGEN (Li et al., 2023). In the domain of subject-driven generation, we benchmark against OmniGen2 (Wu et al., 2025b), DreamO (Mou et al., 2025), UNO (Wu et al., 2025c), XVerse (Chen et al., 2025a), and MIP-Adapter (Huang et al., 2024) to specifically assess identity preservation. For a cutting-edge benchmark, we highlight the latest proprietary models, including Nano Banana (formally named Gemini 2.5 Flash Image, Google’s latest multimodal model) (DeepMind, 2025) and GPT-4o-Image (OpenAI, ), as well as leading open-source models like Qwen-Image-Edit (Wu et al., 2025a) and FLUX.1-Kontext (Labs et al., 2025).
@@ 4.3 Comparison
Table 2: Quantitative results on COCO-MIG and LayoutSAM-Eval. Image-guided methods () use our pre-generated images by FLUX.1-Dev (Labs, 2024b).
@@ 4.4 Ablation Study
Prior work (Zhou et al., 2025b) has demonstrated that MID-19 blocks have the most significant influence on instance-specific attributes. In alignment with this finding, our experiments confirm that applying the ICA mechanism selectively to the MID-19 blocks yields the highest average score of 64.66 and the best IDS score of 32.72. Furthermore, the results in Tab. 3 suggest some potential functional focus across the FLUX-DiT layers, with a detailed analysis provided in Sec. B.6.
@@
To mitigate the model’s tendency to rigidly copy layout images while neglecting instance adaptation (e.g., posture, lighting), we employ Direct Preference Optimization (DPO) (Rafailov et al., 2024), utilizing target images as preferred samples and layout images as less preferred. We first conduct an ablation study on LoRA Rank 256 to determine the optimal $\beta$ coefficient (results summarized in Tab. 4). The optimal $\beta$ (1000) was subsequently applied to our final DPO fine-tuning, which uses the higher capacity LoRA Rank 512. The qualitative validation for the final model is provided in Sec. B.3.3.
@@ 5 Conclusion
Acknowledgments
This work was supported by National Natural Science Foundation of China (U2336212) and the Fundamental Research Funds for the Central Universities (226-202500080).
@@ A.1 Base Model Selection
In our framework, different diffusion backbone models exhibit variations under fine-tuning-free settings for achieving high-fidelity multi-instance generation. We evaluated three variants of the FLUX family of models as potential backbones: FLUX.1-Dev (a general image generation model) (Labs, 2024b), FLUX.1-Fill (a local inpainting model) (Labs, 2024a), and FLUX.1-Kontext (an editing model) (Labs et al., 2025). While existing multi-subject-driven generation methods without layout control (Wu et al., 2025c; Hu et al., 2025) without attention masks have successfully utilized FLUX.1-Dev, our experiments showed a significant limitation: without additional fine-tuning, FLUX.1-Dev failed to produce coherent images when an attention mask was applied. In contrast, both FLUX.1-Fill and FLUX.1-Kontext demonstrated the ability to generate images correctly with the attention mask. Among these two, FLUX.1-Kontext exhibited a noticeably superior capacity for identity preservation. Therefore, we chose FLUX.1-Kontext as the backbone, as it fits better with our framework’s attention-masking strategies and identity preservation goals. To isolate the impact of the backbone, we re-implemented a baseline method MS-Diffusion on FLUX.1-Kontext backbone for ablation studies in Sec. B.7. And we conducted qualitative comparisons with other state-of-the-art methods built upon equivalent FLUX-Family backbones in Sec. D.1.
@@ A.2 Details of Compositing Layout Image
Our Contextual Layout Anchoring (CLA) mechanism relies on a meticulously constructed composite layout image to achieve robust spatial control. The design of this mechanism is conceptually inspired by the principles of multi-feature learning (Yang et al., 2013), which suggests that combining evidence from heterogeneous feature representations can lead to more robust semantic understanding. This process involves two key steps: determining the optimal composition order for all instances and then precisely placing each instance onto the canvas.
A correct composition order is crucial for multi-instance synthesis, especially when handling occlusions and complex overlaps. We propose a dynamic sorting algorithm, Instance Layering Prioritization, which first handles explicit containment relationships by prioritizing instances whose masks are completely contained within another’s. For all other candidate instances, we use a hybrid priority scoring system to simulate the natural layering of objects. We utilize a pre-processing step to obtain each instance’s precise effective area (Ravi et al., 2024). The priority score $P_{i}$ is calculated as:
$P_{i}=\alpha\cdot\mathcal{A}(\text{instance}_{i})+\beta\cdot\left(1-\sum_{j\neq i}\text{IoU}(\text{instance}_{i},\text{instance}_{j})\right)+\lambda\cdot\text{RandomFactor},$
@@ B.1 Experimental Details on LAMICBench++
As shown in Tab. 1, we benchmark our method against several strong baselines, including single-image-editing models (Wu et al., 2025a; Labs et al., 2025) and closed-source commercial models (OpenAI, 2024; DeepMind, 2025). Since these two categories of models require different input modalities and instruction methods, we prepared distinct inputs for a fair evaluation:
@@ B.2 Ablation on Position Indexing
As specified in Sec. 3.2, we extend the position indexing strategy from prior work (Wu et al., 2025c), which uses non-overlapping indices for explicit spatial delineation of objects, crucial for multi-instance distinction. The ablation results on LAMICBench++ are presented in Tab. 5. Our findings validate the importance of the positional encoding components:
@@ B.6 Analysis on Layer-wise ICA Ablation
  • F-19: Strategy to Minimize Modality Interference. Prior work CreatiLayout (Zhang et al., 2024) extending the F-19 blocks has shown a powerful control in both spatial and attribute. And as stated this work, no further modification is made to the other blocks (Single Stream Blocks, i.e., Mid-19 and Back-19 blocks) in order to reduce the potential interference among modalities (i.e., text, layout, and image). It does make sense in considering the Front-19 blocks contains a huge amount of parameters compared to the others. Adding ICA to these blocks is sufficient, while applying ICA to more blocks lead to some side effect like identity distortions. This aligns with our observation that $F+M+B$ and $F+B$ strategies underperform others.
  • M-19: Critical Role in Attribute Binding. Prior work DreamRenderer (Zhou et al., 2025b) demonstrates through ablation study that Mid-19 blocks play a relatively more critical role in attribute binding. Our ablation study validates the finding, for applying ICA to Mid-19 blocks yields the best performance in our task on benchmark score.
@@ B.7 MS-FLUX: A Controlled Case Study of Strategy Failure on Strong Backbone
Previous image generation methods based on older diffusion models (e.g., Stable Diffusion) have shown considerable improvements over their base architectures. For instance, MS-Diffusion (Wang et al., 2025), built on SDXL, achieved image-guided layout control via its Grounding Resampler mechanism based on SD-Unet cross attention architecture. This Resampler fuses reference image content with explicit spatial and phrase information into multi-modal prompt tokens to guide generation.
@@
  • Resampler Initialization and Input Preparation. The Grounding Resampler is initialized within custom LightningModule to match the dimensions of the FLUX Transformer. During the training step, image, layout, and phrase information are prepared.
@@ C.1 Brief Survey and Discussion on Existing Multi-Instance Datasets
Various approaches have been adopted to construct multi-instance datasets. Prior works like OmniGen series (Wu et al., 2025b; Xiao et al., 2024b) and MS-Diffusion (Wang et al., 2025) use a cross-verification method to match and pair group images and corresponding individual images from web images or video frames. And some methods begin to leverage synthetic data for training. For example, UNO (Wu et al., 2025c) employs a co-evolutionary data generation and training process using the model itself, and XVerse (Chen et al., 2025a) add data generated by FLUX.1-Dev as supplementary to high-aesthetic-quality images.
@@ C.2 More Details of IMIG-100K Dataset
We have designed multiple prompt generation templates to include more scenes and styles. These templates are randomly combined and fed to LLMs maximize the diversity of the resulting image generation prompts. We also use different LLMs (i.e. DeepSeek v3 (DeepSeek-AI, 2025), Gemini 2.5 Flash (DeepMind, 2025) and GPT-4o (OpenAI, 2024)) for lexical and stylistic variations. Besides the realistic style (natural style), as illustrated in Fig. 12, we also include extra styles like fantasy and anime, which account for more user scenarios. The proportion of natural scenes remains high, accounting for approximately 85% to 90% of the total data.
@@
Every instance’s is meticulously annotated with detailed spatial information, including the instance phrase, its valid area’s bounding box and mask, and its corresponding position within the ground-truth image. The third subset of IMIG-100K—the flexible composite part—features more detailed annotations. Recognizing that many user scenarios only provide a facial avatar instead of a half or full-body image, we use face-related tools (Xin, 2022; Guo et al., 2021) to restore and standardize the faces in reference images into aligned facial images, as shown in Fig. 14. Face pairs in the ground-truth images (composited images) and corresponding reference images are strictly filtered by face quality and similarity. Crucially, the positions of valid faces are annotated (indicated by the green round bounding boxes in Fig. 14) across composited images, corresponding reference images and aligned face images. This enables a precise position and scale alignment when positioning and sizing reference instances in canvas in training process, instead of relying on a very rough bounding box of the entire instance.
@@
Our dataset has undergone strict quality assessment (Kirstain et al., 2023; Boutros et al., 2023) for both reference images and ground-truth images. And we leverage MLLMs to ensure the semantic consistency between reference images and generated captions. This aligns with recent efforts in multi-context understanding (Xu et al., 2025) and multimodal condition alignment (Zhou et al., 2025a), which both emphasize maintaining logical and semantic integrity across complex visual inputs. Additional samples illustrating the quality and annotation detail are provided in Figs. 13 and 14.
@@ D.2 More Results on COCO-MIG
Full quantitative result on COCO-MIG benchmark is shown in Tab. 8. Our method establishes a new state-of-the-art on all key metrics, achieving the highest average Success Rate (33.12%) and mIoU (65.12%) among all compared methods. Notably, our performance advantage is most pronounced in complex, high-instance-count scenarios. For $L_{4}$, $L_{5}$, and $L_{6}$ levels, our method significantly outperforms all baselines with a Success Rate of 28.12%, 23.12%, and 24.38%, respectively. This demonstrates the robust scalability of our hierarchical architecture to maintain both layout and identity control in intricate scenes. While some competitors show slightly higher scores on individual metrics like Global Clip (CreatiLayout (Wu et al., 2025c)) or Local Clip (InstanceDiffusion (Wang et al., 2024)), our approach achieves a superior overall balance. The consistently high mIoU scores across all complexity levels and our leading Instance Success Rate on the most challenging cases further validate our model’s ability to master both precise instance placement and high-fidelity attribute preservation.
More qualitative results on COCO-MIG are shown in Fig. 17. Our method demonstrates a clear qualitative advantage over existing models on the COCO-MIG benchmark. In the first example, other methods fail to correctly generate the blue vase, with issues ranging from incorrect position to instance merging. Our model, in contrast, precisely renders the vase as intended. Similarly, for the ”green potted plant”, our approach correctly applies the “green” attribute to the pot, whereas competitors fail to do so. This highlights our superior ability to handle holistic subject identity. Furthermore, while some baselines correctly generate the requested objects in the third and fourth examples, their outputs often lack aesthetic harmony and visual coherence. Our method consistently produces images that are not only accurate but also visually pleasing and well-composed. These results underscore two key advantages of our framework: (1) Compared to other image-guided methods like MS-Diffusion (Wang et al., 2025), our approach offers significant superiority in layout control (2) Our dedicated identity preservation mechanism provides more robust and reliable subject fidelity than the attribute-based control of text-guided methods, particularly in intricate, multi-instance scenes.
@@ Appendix E Limitations and Future Work
While our framework demonstrates state-of-the-art performance in multi-instance generation, it is not without limitations. A primary challenge stems from our model’s strong emphasis on identity preservation. When inconsistencies exist between the provided reference images or between the images and the text prompt, our model tends to prioritize maintaining the identities of the reference subjects. This can sometimes lead to a lack of flexibility in adjusting attributes such as lighting, color, or pose, which may compromise the overall visual harmony and text-image consistency of the final output. This trade-off between identity fidelity and contextual flexibility represents an important area for future research. In the future, we plan to explore more dynamic attention mechanisms that can better balance these competing demands, allowing for more flexible style and attribute transfer. Drawing on zero-shot consistency priors (Zhang et al., 2025b), we also aim to incorporate broader generative knowledge to improve the overall harmony of multi-instance scenes. Furthermore, while ContextGen focuses on 2D image synthesis, extending identity-consistent control to 3D assets remains a promising direction (Xu et al., 2024; Shen et al., 2024).